# Mechanically Programming the Cross-Sectional Shape of Soft Growing Robotic Structures for Patient Transfer

### v1

@@ V CONCLUSION
This paper presents a method of mechanically programming the cross-sectional shape of soft everting robotic structures using flexible strips that constrain the radial expansion between points along the outer membrane. This method is the first to enable these capabilities while maintaining the full multi-axis bending flexibility of the membrane when deflated. We also present methods for fabricating and implementing these structures for robotic slings, develop and validate an analytical model relating its design specifications and fabrication parameters, and prototype a full soft growing robotic sling system to demonstrate its use for assisting a single caregiver in harnessing and transferring a patient from a bed to a chair. Unlike traditional slings, our prototype places itself under the patient automatically via eversion, eliminating the need for the caregiver to do so manually. Due to its multi-axis bending compliance once deflated, it could then subsequently be used exactly like a standard sling to perform the lift and transfer. Our demonstration is the first to harness, lift, and transfer a patient using a soft growing robot with a sling-like flattened cross-section that distributes contact pressures over a wide area. Future work will include the integration of bend mechanisms to enable eversion/retraction through bent pathways (e.g. under a patient sitting on a chair), further generalization of the geometric model to account for non-symmetrical designs, and more in-depth experimental characterization.

### v3

@@ V CONCLUSION
This paper presents a method of mechanically programming the cross-sectional shape of soft everting robotic structures using flexible strips that constrain the radial expansion between points along the outer membrane. This method is the first to enable these capabilities while maintaining the full multi-axis bending flexibility of the membrane when deflated. We also present methods for fabricating and implementing these structures for robotic slings, develop and validate an analytical model relating its design specifications and fabrication parameters, and prototype a full soft growing robotic sling system to demonstrate its use for assisting a single caregiver in harnessing and transferring a patient from a bed to a chair. Unlike traditional slings, our prototype places itself under the patient automatically via eversion, eliminating the need for the caregiver to do so manually. Due to its multi-axis bending compliance once deflated, it could then subsequently be used exactly like a standard sling to perform the lift and transfer. Our demonstration is the first to harness, lift, and transfer a patient using a soft growing robot with a sling-like flattened cross-section that distributes contact pressures over a wide area. Future work will also include the integration of bend mechanisms to enable eversion/retraction through bent pathways (e.g. under a patient sitting on a chair), and further generalization of the geometric model to account for non-symmetrical designs. There is also room for more in-depth modeling and experimental characterization of our system under different external loads during growth and retraction, as well as further user studies with larger and more diverse participant sample sizes. Although the participants did not need to lift or support their upper body to aid deployment, a future study could include soft contact sensing to better understand the safety and adaptability of our prototype to scenarios with less responsive users.